Edelman and Gronkowski want to play a game of beach volleyball, if the beach ball has a volume of 400 L at a temperature of 500 k, what will the temperature of the balloon be if the volume decreases to 176 L?


1
Expert's answer
2021-06-14T13:36:09-0400

V1 = 400L , T1 = 500K , V2 = 176L , T2 = ?


"\frac{V_1}{T_1} = \frac{P_2}{T_2}"


"\frac{400}{500} = \frac{176}{T_2}"


T2 = "\frac{500×176}{400}"


T2 = 220K